the resonant frequency F of a closed air column is known to depend on the pressure and the density, p of the air and the length L of the air column. how are these quantities related?
1
Expert's answer
2019-05-23T09:44:00-0400

The resonant frequency F of a closed air column:


"F=\frac{v_{sound}}{4L}"

At the same time:


"v_{sound}=\sqrt{\frac{K_s}{\rho}}"

Consequently:


"F=\frac{1}{4L}\sqrt{\frac{K_s}{\rho}}"

Where "K_s" is a coefficient of stiffness, "\rho" -air density, "L" -length of the air column.